OpenStudy (anonymous):

Find all numbers for which the rational espression is undefined. q^3 - 7q/ q^2-4

OpenStudy (anonymous):

all real numbers except for 2 and -2

OpenStudy (anonymous):

Q = ?

OpenStudy (anonymous):

q = 2 or q = -2 gives zero denominator - sorry - i misread the question. expression is undefined for these 2 values of q

OpenStudy (anonymous):

q^30 7q/q^2 - 4

OpenStudy (anonymous):

Oh crap...q^3 - 7q / q^2 -4

OpenStudy (anonymous):

i've assumed this is q^3 - 7q -------- q^2 -4 is this right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok - well thats the answer :- q = 2,-2 gives an undefined value for the expression because (2)^2 - 4 = 0 and (-2)^2 - 4 = 0 and something / 0 is undefined

OpenStudy (anonymous):

That was a good answer. Just has a baby night (nephew) so my mind is not right.

OpenStudy (anonymous):

good rhyming there :)